OpenStudy (arbyj):

helper will get a medal !! Write the standard form of the line that passes through the point (-2, 4) and is parallel to x - 2y = 6

OpenStudy (pphalke):

Do you know standard form ?

OpenStudy (arbyj):

yes ax+by=c

OpenStudy (pphalke):

Have you maybe tried plugging in the numbers

OpenStudy (arbyj):

plugging in the points?

OpenStudy (lapinblanc):

Plugging in the points just gives you a number. That won't help you. The purpose of the equation provided is to show you the slope of the parallel line. Parallel lines have the EXACT same slope. So, -rearrange the given equation, identify the slope (it is the number that ends up in front for the x) -use this formula \[y-y _{1}=m(x-x _{1})\] Where y sub 1 and x sub 1 are your provided points (-2, 4) -Rearrange into standard form\[y=mx+b\]

OpenStudy (lapinblanc):

And m is your slope.

OpenStudy (photon336):

yeah do you know like conceptually what parallel means? it means that the two lines never intersect which is that they must have the same slope.

OpenStudy (arbyj):

yes, but it's just the equation im having trouble finding

OpenStudy (lapinblanc):

Did the above information give you what you needed or are you having trouble rearranging it?

OpenStudy (arbyj):

im having trouble reearranging

OpenStudy (lapinblanc):

I will use the provided equation to show you an example: x-2y=6 Since we need the equation to be form y=mx+b We subtract x from both sides x(-x)-2y=6-x x on the left cancels out -2y=6-x we need y alone, so we divide both sides by -2 (or multiply by the reciprocal, same thing) -2y/-2=(6-x)/-2 the -2 on the left cancels out bc anything divided by itself is 1 y=(6-x)/-2 But you want the slope so you only care about the number in front of the x, so you will want to break it up. \[\frac{ 6-x }{ -2 }\] \[\frac{ 6 }{ -2 } - \frac{ x }{ -2 }\] which is the same as \[-3+\frac{ -x }{ -2 }\] cancel out the negatives and treat the remaining number as a coefficient \[-3+\frac{ 1 }{ 2 }x\] Your slope is the number that remains in front of the x.

OpenStudy (lapinblanc):

All those equation have y= I got lazy, sorry.

OpenStudy (arbyj):

its okay haha i got it now

OpenStudy (arbyj):

thank you!

OpenStudy (lapinblanc):

^__^ Excellent

OpenStudy (mathmate):

Alternative way without finding slope: If we need a line parallel to another given in standard form Ax+By=C, and the new line has to pass through a given point P(x0,y0). The new line in standard form is given by Ax+By=k where k=A(x0)+B(y0), in other words, the new line is \(Ax+By=A(x0)+B(x0)\) Say given a line L1: x-2y=6, need a line L2 parallel to L1 and through (-2,4), L2: Ax+By=A(x0) +B(y0), where (x0,y0)=(-2,4) x-2y=1(-2)-2(-4)=-10 So the required line is L2: x-2y=-10